OpenStudy (anonymous):

Given the function f described by f(x)=x+6,find each of the following f(-18)=

OpenStudy (anonymous):

-12

OpenStudy (angela210793):

f(-18)=-18+6=-12 all u have to do is replace -18 instead of x

OpenStudy (anonymous):

f(-2)=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

f(6)=

OpenStudy (angela210793):

same as i told u...replace the no. instead of x

OpenStudy (anonymous):

12

OpenStudy (anonymous):

f(d+18)=

OpenStudy (anonymous):

d+24

OpenStudy (anonymous):

Come on just plug in the value of x in the expression